OpenStudy (anonymous):

Find the roots of (5x + 6)(3x + 2)(4x - 4)

OpenStudy (anonymous):

@mathmale

hero (hero):

So P(x) = (5x + 6)(3x + 2)(4x - 4)

hero (hero):

Set P(x) = 0: (5x + 6)(3x + 2)(4x - 4) = 0

hero (hero):

Then by zero product property: 5x + 6 = 0 3x + 2 = 0 4x - 4 = 0 Solve for x for each

OpenStudy (anonymous):

is it D?

hero (hero):

Do you know how to solve an equation for x? Also, avoid just saying the answer is D. If you believe you have the correct answer, demonstrate why your choice is correct. Can you show the work you did to arrive at your result?